# Cuipo Jiang and Jingtian Song

### v2

@@ 3.1 Kashiwara-Tanisaki Characters
It is obvious that $-\Lambda_{0},k\Lambda_{0}\in{\mathcal{C}}^{+}$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$. Notice that for any $w\in\widehat{W}(-\Lambda_{0})$ which is the longest element of $w\widehat{W}_{0}(-\Lambda_{0})$, $\sigma(w)$ is the longest element of $\sigma(w)\widehat{W}_{0}(k\Lambda_{0})$. Then by Theorem 3.2 and (3.9)-(3.11) ( see also Theorem 11 of [Fie06] from the category point of view),
${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(\sigma(y)\circ(k\Lambda_{0}))),$ (3.12)
which is equivalent to
$\widehat{R}{\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1)e^{\sigma(y)\circ k\Lambda_{0}}.$ (3.13)
Then by Lemma 3.3, Lemma 3.4, and Theorem 3.5,
$\widehat{R}{\rm ch}L(k\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
Furthermore, (3.12) together with Lemma 3.6 means that $L(y\circ(-\Lambda_{0}))$ for some $y\in\widehat{W}(-\Lambda_{0})$ is isomorphic to a subquotient of $M(w\circ(-\Lambda_{0}))$ if and only if $L(\sigma(y)\circ k\Lambda_{0})$ is a subquotient of $M(\sigma(w)\circ k\Lambda_{0})$.
Case 1   $n=3$, $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$. Let $u^{1}$ and $u^{2}$ be the singular vectors given in (1) of Theorem 3.7. The weights of $u^{1}$ and $u^{2}$ with respect to $\widehat{\mathfrak{h}}$ are $-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2}$ and $-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}$, respectively. It is easily checked that
$s_{\alpha_{2}}\circ(t_{\alpha_{1}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2},$ (3.14)
@@
$s_{\alpha_{1}}\circ(t_{\alpha_{2}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}.$ (3.15)
Thus $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ of wights $s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$ and $s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$, respectively. It is easy to check that
$s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2},$ (3.16)
@@
$s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2},$ (3.17)
By (3) of Theorem 3.7, the ideal generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$. It follows that the ideal generated by $v^{1}$ and $v^{2}$ is the maximal ideal of $V^{k}(sl_{3})$. We complete the proof of (1).
Case 2  $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$. Let $u$ be the singular vector in (2) of Theorem 3.7. The weight of $u$ with respect to the Cartan subalgebra $\widehat{\mathfrak{h}}$ is $-\Lambda_{0}-2\delta+\theta+\beta$, where $\beta=\alpha_{2}+\cdots+\alpha_{n-2}$. It is easily checked that
@@
Then $V^{k}(sl_{n})$ has a singular vector $w^{1}$ of weight $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-q\beta}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$. It can be checked directly that
@@
By (4) of Theorem 3.7, the ideal generated by $u$ is the maximal ideal of $V^{-1}(sl_{n})$. It follows that the ideal generated by $w^{1}$ is the maximal ideal of $V^{k}(sl_{n})$. ∎
Remark 3.9.
The method here can be used to the subregular case for $\mathfrak{g}$ of other types.
@@ 4.1 The associated variety of $L_{-1}(sl_{3})$
are singular vectors of $V^{-1}(sl_{3})$. Then we have
@@
$\begin{array}[]{ll}x\equiv&-e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)h_{2}(-1){\bf 1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1){\bf 1}\\ \\ &-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)(2h_{1}+h_{2})(-1){\bf 1}\\ \\ &+2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{2}}(-1){\bf 1}({\rm mod}~{}F^{1}(V^{-1}(sl_{3})))\\ \\ y\equiv&h_{1}(-1)h_{2}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{3}}(-1){\bf 1}\\ \\ &-3e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)f_{\alpha_{2}}(-1){\bf 1}-e_{\alpha_{1}}(-1)f_{\alpha_{1}}(-1)h_{1}(-1){\bf 1}\\ \\ &+e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}\\ \\ &-e_{\alpha_{2}}(-1)f_{\alpha_{2}}(-1)h_{2}(-1){\bf 1}\ ({\rm mod}~{}F^{1}(V^{-1}(sl_{3}))).\end{array}$
@@
Then by Lemma 4.1, $p_{i}\in\overline{{\mathcal{I}}_{-1}}$, $1\leqslant i\leqslant 4$.
@@
This implies that $G\cdot(f_{\alpha_{1}}+f_{\alpha_{2}})$ and $G\cdot({\mathbb{C}}^{*}\lambda+f_{\theta})$ could not belong in $X_{L_{-1}(sl_{3})}$. Furthermore, for regular semisimple vector $h$, all $(h_{1}|h),\ (h|h_{2}),(h|h_{1}+h_{2})$ are non-zero. So

### v3

@@ 3.1 Kashiwara-Tanisaki Characters
$l_{-\Lambda_{0}}(y)=l_{k\Lambda_{0}}\sigma(y),\ l_{-\Lambda_{0}}(w)=l_{k\Lambda_{0}}\sigma(w).$ (3.13)
Let $w\in\widehat{W}(-\Lambda_{0})$ be such that $w$ is the longest element of $w\widehat{W}_{0}(-\Lambda_{0})$. Then by (3.11) and (3.13), $\sigma(w)$ is the longest element of $\sigma(w)\widehat{W}_{0}(k\Lambda_{0})$. By Theorem 3.3,
${\rm ch}(L(w\circ(-\Lambda_{0})))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(y\circ(-\Lambda_{0}))),$ (3.14)
and
${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{\sigma(w)\leqslant_{k\Lambda_{0}}\sigma(y)\in\widehat{W}(k\Lambda_{0})}(-1)^{l_{k\Lambda_{0}}(\sigma(y))-l_{k\Lambda_{0}}(\sigma(w))}Q^{k\Lambda_{0}}_{\sigma(w),\sigma(y)}(1){\rm ch}(M(\sigma(y)\circ k\Lambda_{0})).$ (3.15)
Since $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ via the automorphism $\sigma$, by the definition of $P^{\lambda}_{y,w}(q)$ and $Q^{\lambda}_{y,w}(q)$ [KL79], for $y,w\in\widehat{W}(-\Lambda_{0})$,
@@
Then together with (3.12)-(3.13), we have
${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(\sigma(y)\circ k\Lambda_{0})).$ (3.17)
The equation systems (3.14) and (3.17) also implies that for $w,y\in\widehat{W}$,
$[M(w\circ(-\Lambda_{0})):L(y\circ(-\Lambda_{0}))]=[M(\sigma(w)\circ k\Lambda_{0}):L(\sigma(y)\circ k\Lambda_{0})],$ (3.18)
where $[M:L(\mu)]$ denotes the multiplicity of the irreducible subquotient $L(\mu)$ in $M$, for a module $M$ in category ${\cal O}$.
By the character formulas given in [KK79] for generalized Verma modules of affine Lie algebras induced from integrable modules of the associated simple Lie algebras, one can easily deduce that
$[V(w\circ(-\Lambda_{0})):L(y\circ(-\Lambda_{0}))]=[V(\sigma(w)\circ k\Lambda_{0}):L(\sigma(y)\circ k\Lambda_{0})],$ (3.19)
where $V(w\circ(-\Lambda_{0}))$ is the generalized Verma module of the affine Lie algebra $\widehat{sl_{n}}$ induced from the finite-dimensional $sl_{n}$-module $L(\overline{w\circ(-\Lambda_{0}}))$, and $V(\sigma(w)\circ(k\Lambda_{0}))$ is the generalized Verma module induced from $L(\overline{\sigma(w)\circ k\Lambda_{0}})$.
Notice that (3.17) can also be described as
$\widehat{R}{\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1)e^{\sigma(y)\circ k\Lambda_{0}},$ (3.20)
where
$\widehat{R}=\prod_{\alpha\in\widehat{\Delta}_{+}}(1-e^{-\alpha})^{{\rm dim}{\widehat{\mathfrak{g}}}_{\alpha}},\ \mathfrak{g}=sl_{n}.$
By (3.9),
$\widehat{R}{\rm ch}L_{-1}(sl_{n}))=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
By Lemma 3.7 and Lemma 3.8, for $w\in W$, $\gamma\in Q$,
$\sigma(wt_{\gamma})=wt_{q\gamma}.$
Then we have
(3.10) is proved.
Set
$\Lambda^{(-1)}=\widehat{W}(-\Lambda_{0})\circ(-\Lambda_{0})=\{w\circ(-\Lambda_{0})|w\in\widehat{W}(-\Lambda_{0})\}$
@@
$\Lambda^{(k)}=\widehat{W}(k\Lambda_{0})\circ k\Lambda_{0}=\{w\circ(k\Lambda_{0})|w\in\widehat{W}(k\Lambda_{0})\}.$
Let ${\cal O}_{\Lambda^{(-1)}}$ and ${\cal O}_{\Lambda^{(k)}}$ be the corresponding indecomposable blocks, respectively. Then by Theorem 3.5 and Lemma 3.7,
${\cal O}_{\Lambda^{(-1)}}\cong{\cal O}_{\Lambda^{(k)}}.$ (3.21)
We need to explain the equivalence (3.21) briefly. Notice that both $-\Lambda_{0}$ and $k\Lambda_{0}$ are dominant. Then for $w\in\widehat{W}(-\Lambda_{0})$, $L(w\circ(-\Lambda_{0}))$ (resp. $L(\sigma(w)\circ k\Lambda_{0})$) has a projective cover $P(w\circ(-\Lambda_{0}))$ (resp. $P(\sigma(w)\circ k\Lambda)$) in ${\cal O}_{\Lambda^{(-1)}}$ (resp. ${\cal O}_{\Lambda^{(k)}}$). Set
${\cal P}_{\Lambda^{(-1)}}=\{P(w\circ(-\Lambda_{0}))|w\in\widehat{W}(-\Lambda_{0})\}$
@@
${\cal P}_{\Lambda^{(k)}}=\{P(\sigma(w)\circ k\Lambda_{0})|w\in\widehat{W}(-\Lambda_{0})\}.$
${\cal P}_{\Lambda^{(-1)}}$ (resp. ${\cal P}_{\Lambda^{(k)}}$) is a faithful set of small projectives in the sense of [Mit72]. View ${\cal P}_{\Lambda^{(-1)}}$ (resp. ${\cal P}_{\Lambda^{(k)}}$) as a full subcategory of ${\cal O}_{\Lambda^{(-1)}}$(resp. ${\cal O}_{\Lambda^{(k)}}$). Let ${\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(-1)}}^{\rm opp}}$ (resp. ${\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(k)}}^{\rm opp}}$) be the category of additive functors ${\cal P}_{\Lambda^{(-1)}}\rightarrow{{\mathbb{C}}}-{\rm mod}$ (resp. ${\cal P}_{\Lambda^{(k)}}\rightarrow{{\mathbb{C}}}-{\rm mod}$). By Theorem 3.1 of [Mit72],
${\cal O}_{\Lambda^{(-1)}}\rightarrow{\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(-1)}}^{\rm opp}}\ \ ({\rm resp.}\ {\cal O}_{\Lambda^{(k)}}\rightarrow{\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(k)}}^{\rm opp}})$
$M\rightarrow{\rm Hom}(\cdot,\ M)\ \ \ ({\rm resp.}\ \tilde{M}\rightarrow{\rm Hom}(\cdot,\ \tilde{M})).$
is an equivalence of ${{\mathbb{C}}}$-categories. Theorem 3.5 (i.e., Theorem 11 of [Fie06]) tells us that the Coxeter group automorphism $\sigma$ determines an equivalence between the categories ${\cal O}_{\Lambda^{(-1)}}$ and ${\cal O}_{\Lambda^{(k)}}$ via the above ${\mathbb{C}}$-category equivalences (c.f. [Fie06] for details). Let $F$ be the covariant functor from ${\cal O}_{\Lambda^{(-1)}}$ to ${\cal O}_{\Lambda^{(k)}}$. Then in particular, $F$ maps the class of Verma module $M(w\circ(-\Lambda_{0}))$ to the class of Verma module $M(\sigma(w)\circ k\Lambda_{0})$, the class of simple module $L(w\circ(-\Lambda_{0}))$ to the class of simple module $L(\sigma(w)\circ k\Lambda_{0})$, and it is easy to see that $F$ maps the class of $V(w\circ(-\Lambda_{0}))$ to the class of $V(\sigma(w)\circ k\Lambda_{0})$, for $w\in\widehat{W}(-\Lambda_{0})$.
Case 1   $n=3$, $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$. Let $u^{1}$ and $u^{2}$ be the singular vectors given in (1) of Theorem 3.10 with highest weights $-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2}$ and $-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}$, respectively. It is easy to check that
$s_{\alpha_{2}}\circ t_{\alpha_{1}}\circ(-\Lambda_{0})=s_{\alpha_{2}}\circ(t_{\alpha_{1}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2},$ (3.22)
and
$s_{\alpha_{1}}\circ t_{\alpha_{2}}\circ(-\Lambda_{0})=s_{\alpha_{1}}\circ(t_{\alpha_{2}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}.$ (3.23)
Then by (3.18) and Lemma 3.4, $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ of weights $s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0}$ and $s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0}$, respectively. It is easy to check that
$s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0}=s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2},$ (3.24)
and
$s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0}=s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2}.$ (3.25)
Since
$[V^{-1}(sl_{3}):L(s_{\alpha_{1}}t_{\alpha_{2}}\circ(-\Lambda_{0}))]=[V^{-1}(sl_{3}):L(s_{\alpha_{2}}t_{\alpha_{1}}\circ(-\Lambda_{0}))]=1,$
by (3.19),
$[V^{k}(sl_{3}):L(s_{\alpha_{1}}t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0})]=[V^{k}(sl_{3}):L(s_{\alpha_{2}}t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0})]=1.$ (3.26)
Let $\tilde{I}$ be the ideal of $V^{k}(sl_{3})$ generated by $v^{1}$ and $v^{2}$. Since $I$ is the maximal ideal of $V^{-1}(sl_{3})$, we have the following exact sequence:
$0\rightarrow I\rightarrow V^{-1}(sl_{3})\rightarrow L_{-1}(sl_{3})\rightarrow 0$
Then we have the exact sequence:
$0\rightarrow F(I)\rightarrow V^{k}(sl_{3})\rightarrow L_{k}(sl_{3})\rightarrow 0,$
since $F(V^{-1}(sl_{3}))=V^{k}(sl_{3})$, $F(L_{-1}(sl_{3}))=L_{k}(sl_{3})$.
On the other hand, since $I$ is a quotient of $M(s_{\alpha_{1}}\circ t_{\alpha_{2}}\circ(-\Lambda_{0}))\oplus M(s_{\alpha_{2}}\circ t_{\alpha_{1}}\circ(-\Lambda_{0}))$, it follows that $F(I)$ is a quoitent of $M(s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0})\oplus M(s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0})$. This together with (3.26) means that $F(I)$ must be the ideal $\tilde{I}$ generated by $v^{1}$ and $v^{2}$. Since $F(I)$ is maximal, it follows that $\tilde{I}$ is maximal.
Case 2  $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$. Let $u$ be the singular vector in (2) of Theorem 3.10. The weight of $u$ with respect to the Cartan subalgebra $\widehat{\mathfrak{h}}$ is $-\Lambda_{0}-2\delta+\theta+\beta$, where $\beta=\alpha_{2}+\cdots+\alpha_{n-2}$. It is easy to check that
@@
Then by (3.18) and Lemma 3.4, $V^{k}(sl_{n})$ has a singular vector $w^{1}$ of weight $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0}$. One can easily check that
@@
By (4) of Theorem 3.10, the ideal $J$ generated by $u$ is the maximal ideal of $V^{-1}(sl_{n})$. Then we have the exact sequence:
$0\rightarrow J\rightarrow V^{-1}(sl_{n})\rightarrow L_{-1}(sl_{n})\rightarrow 0$ (3.27)
By the fact that $F(V^{-1}(sl_{n}))=V^{k}(sl_{n})$, and $F(L_{-1}(sl_{n}))=L_{k}(sl_{n})$, we have
$0\rightarrow F(J)\rightarrow V^{k}(sl_{n})\rightarrow L_{k}(sl_{n})\rightarrow 0$ (3.28)
Notice that $J$ is a quotient of $M(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-\beta})\circ(-\Lambda_{0}))$, it follows that $F(J)$ is a quotient of $M((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0})$. Then $F(J)$ is generated by $w^{1}$, since
$[V^{-1}(sl_{n}):L((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-\beta})\circ(-\Lambda_{0}))]=[V^{k}(sl_{n}):L((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0})]=1.$
By (3.28), $F(J)$ is maximal. We complete the proof of (2). We remark that (3.28) can also be deduced from (3.27) and (3.19) without using the category equivalence (3.21). ∎
@@ 4.1 The associated variety of $L_{-1}(sl_{3})$
are singular vectors of $V^{-1}(sl_{3})$. Then we have the following lemma.
@@
$\begin{array}[]{ll}x\equiv&-e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)h_{2}(-1){\bf 1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1){\bf 1}\\ \\ &-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)(2h_{1}+h_{2})(-1){\bf 1}\\ \\ &+2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{2}}(-1){\bf 1}({\rm mod}~{}F^{1}(V^{-1}(sl_{3})))\\ \\ y\equiv&h_{1}(-1)h_{2}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}\\ \\ &-3e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)f_{\alpha_{2}}(-1){\bf 1}-e_{\alpha_{1}}(-1)f_{\alpha_{1}}(-1)h_{1}(-1){\bf 1}\\ \\ &+e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}\\ \\ &-e_{\alpha_{2}}(-1)f_{\alpha_{2}}(-1)h_{2}(-1){\bf 1}\ ({\rm mod}~{}F^{1}(V^{-1}(sl_{3}))).\end{array}$
@@
Then by Lemma 4.1, $p_{i}\in\overline{{\mathcal{I}}_{-1}}$, $1\leqslant i\leqslant 4$. We have the following result which is also covered in [AFK24] by different method.
@@
This implies that $G.(f_{\alpha_{1}}+f_{\alpha_{2}})$ and $G.({\mathbb{C}}^{*}\lambda+f_{\theta})$ could not belong in $X_{L_{-1}(sl_{3})}$. Furthermore, for regular semisimple vector $h$, all $(h_{1}|h),\ (h|h_{2}),(h|h_{1}+h_{2})$ are non-zero. So